OpenStudy (anonymous):

A square plywood platform has a perimeter that is 6 times the length of a side, decreased by 6. Find the length of a side. (Recall the formula P = 4s.)

OpenStudy (anonymous):

initial perimeter will be p1=4s when side is s now side has decreased to (s-6) now perimeter isp2= 6s so equate both the perimeters p1=p2 4s=6(s=6) 4s=6s−36 2s=36 s=18 so the side will be 18.

OpenStudy (anonymous):

in that not the aanswer this shoulld be P = 4s = 6s-6 6 = 2s s =3

OpenStudy (anonymous):

if you are able to explain your own work, that is awesome

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

:)